OpenStudy (anonymous):

Solve 2x + 5x + x + 3 = 3x + x + 5. Simplify: Show me the steps please

OpenStudy (anonymous):

2x + 5x +x -3x -x = 5 - 3 4x = 2 x=0.5

OpenStudy (anonymous):

Well I know the answer,but I just want to know the steps to solve it.Thanks though.

OpenStudy (anonymous):

You put the x's on one side , the integers on the other side , the solve it You end up with ax=b ..... then x= b/a

OpenStudy (mathstudent55):

Solve showing step-by-step work: 2x + 5x + x + 3 = 3x + x + 5 1. Add like terms together on left side and on right side (2x + 5x + x) + 3 = (3x + x) + 5 8x + 3 = 4x + 5 2. Subtract 4x from both sides to get all x's on left side: 8x + 3 - 4x = 4x + 5 - 4x 8x - 4x + 3 = 4x - 4x + 5 4x + 3 = 5 3. Subtract 3 from both sides to get all numbers on right side 4x + 3 - 3 = 5 - 3 4x = 2 4. Divide both sides by 4 4x/4 = 2/4 5. Simplify both sides: x = 1/2